If f(x) = -7x + 5x2 + 10, what does f(-2) equal?
Answer by amarjeeth123(569)

f(x) = -7x + 5x2 + 10
f(-2)=5(-2)^2-7(-2)+10=5(4)+14+10=20+24=44